OpenStudy (anonymous):

Evaluate 6^(log14/log6)

OpenStudy (anonymous):

any guesses for this one?

OpenStudy (anonymous):

I used log rules but I didn't get the right answer. I tried to go log base of 6 with 0= log14/log6 as well.

OpenStudy (anonymous):

by definition \[b^x=e^{x\log(b)}\] so \[6^{\frac{\log(14)}{\log(6)}}\] \[=e^\frac{\log(14)}{\log(6)\times \log(6)}\] \[=e^{\log(14)}=14\]

OpenStudy (anonymous):

ok i made a typo on line 3 above

OpenStudy (anonymous):

thanks, 14 is the right answer!

OpenStudy (anonymous):

\[e^{\frac{\log(14)}{\log(6)}\log(6)}=e^{\log(14)}=14\]

OpenStudy (anonymous):

yw, sorry about the typo